Text Size:
Smaller Text Normal Text Larger Text


Follow us on Twitter

Changes done to FIES and to related things

This log was started Jun 2007, when FIES had been fully functional as a common-user instrument for about 9 months, in its own building.
  • 18-22 June 2007: installation of fiber unit B.
    This unit consists of 3 fibers only: F1 low-res, F3 med-res, and F4 hi-res.

    The previous bundle (refered to as "fiber unit A" or "the old fiber unit") is considered the spare bundle from now on.

  • 18-22 June 2007: moved bottom calibration unit to front room

  • 18-22 June 2007: added extension fiber to F5 (calibration fiber) of the old fiber unit A, such that it can reach the new location of the bottom calibration unit

  • 18-22 June 2007: installed a flexible tube to lead the CCD-controllers heat out of the spectrograph box. We made a hole in the box for this.

  • 18-22 June 2007: cleaned pickoff mirrors (FIES/STANCAM) with distilled water

  • 27 June 2007:
    installed new ThAr lamp in bottom calib unit, as the old one died.
    The entrance terrace was painted white.
    New virtual ground (earth) installed infront of building entrance (dug-in under mountain surface).
    The west-wall out-side metallic isolation installed.

  • 6 July 2007: new big fan has been installed in front room.

  • 26 July: temperature monitoring sensors re-installed

  • 23 Aug: due to the unstable CCD temperature, it was decided to set the FIES CCD reference temperature from -120 to -105C.

  • 11 Dec 2007: flushed the cryotiger system, and put the CCD reference temperature back to -120.

    Carlos: What has been done about FIES cryotiger is de-contamination of all the system. If anybody want to do this, basically, you need the manual which we have copied from the Liverpool telescope and corrected, all the tools which are also coming from there, a spare PT13 gas bottle, and a full big dry nitrogen bottle.

    The basic is: empty the system from any gas in 2 parts, head and tubes in one side and the compressor in the other side; flush them with dry nitrogen; make vacuum; and refill with gas. If you want more details, have a look to the manual which is in the staff office at orm with the NOTCAM things.

  • 11 Jan 2008: disconnected the front-room cooling-fan mains cable, to prevent it to suck in more dust.

  • 11 Jan 2008: vacuum cleaned FIES room and front room; cleaned top of FIES white box.

  • 11 Jan 2008: found light leak. The new spectrograph-shutter opening is big and was not baffled off. Consequently, light emitted by the electronics (CCD controller and/or shutter board) could enter the spectrograph. Made temporary fix consisting of a slab of black rubber placed over the area of the fiber exit and spectrograph entrance.

  • 11 Apr 2008: cleaned top of black box.

  • 14 Apr 2008: because of continuing problems with the illumination of the low-res fiber by the top calibration unit, the gain-RON monitoring is now using fiber #3 .

  • 11-15 Apr 2008: installed exposure-meter head in spectrograph, which involved opening white and black box, and slight moving of CCD controller.

  • 21 Apr 2008: set a new format for FIES CCD images: 2198x2052. It used to be 2148x2052. There are now 50 extra overscan columns.

  • 23 Apr 2008: cleaned FIES pickoff mirror in adapter, and removed spare bundle from adapter. The spare bundle (A) is now coiled up under the fork.

  • 18 Jun 2008: moved electronics of exposure meter to front room. And replaced a dodgy ribbon cable.

  • [ 19 Jun 2008: first star light on exposure meter. ]

  • 4 Jul 2008: GC installed final components in the focus electronics of FIES. Needs testing.

  • 8 Aug 2008: started-up newly installed front-room cooling

  • 23 Sep 2008: mechanically fixed the top-calib unit, after a low-voltage event caused the arm/mask switches to be blind to mechanical triggers.

  • 23 Sep 2008: removed fiber bundle A from the telescope, to be refurbished into fiber bundle C.

  • 10 Dec 2008: cleaned FIES building, and outside of white box.

  • 13 Jan 2009: realligned the calibration arm with the calibration fiber at the adapter.

  • 27 May 2009: refocussed FIES spectrograph using fiber 1. Changed focus position from 196 to 198. No significant change in PSF of fiber 1.

  • 1 Jun 2009: cleaned STANCAM + FIES pickoff mirrors with alcohol. FIES mirror reflectivity before cleaning: 76-82%. Shows wear, and a large drooling stain.

  • 6 July: pumped cryostat during telescope aluminizing.

  • 14 Aug 2009: dewar was back flushed. See fault 1216.

  • 15 Sept 2009: changed both fans in the FIES CryoTiger compressor. See fault 1216.

  • 28 Sept 2009: removed black cylindrical filter from the CryoTiger hose loop. See fault 1216. This did not solve the problem.

  • 28 Sept 2009: installed fiber bundle C, running from the hole under the telescope to the spectrograph. Fiber bundle B was marked with red tape in the cathedral/cellar.

  • 29 Sept 2009: to flush the system, reversed (swapped) the tubes of the CryoTiger cooling system, for 2 hours, and swapped back again. See fault 1216. Started pumping.

  • 30 Sept 2009: switched on the CryoTiger.

  • 13 Oct 2009: faulty ThAr lamp in bottom calib-unit replaced.

  • 26 Oct 2009: pumped cryo and connected pressure sensor.
    Added 3 clamps to properly hold the CCD-controller duct to the south side of the spectrograph box.

  • 29 Oct 2009: connected back the black cylindrical filter in cryotiger hose loop.

  • 2 Nov 2009: checked if the Xwings of ThAr line are due to the crossdispers' edges, by masking off the edges. The wings are not due to the CD edges.
    Measured the reflectivity of the 2 collimator mirrors in FIES.
    Pumped the cryo again.

  • 3 Nov 2009: Switched off pumping, reverse flushed the warm dewar first thing in the morning, swapped to the new filter, and added some gas and cool.

  • 5 Nov 2009: swapped to fiber bundle C for its first light.

  • 6 Nov 2009: swapped back to fiber bundle B.

  • 16 Nov 2009: installed an XY table at the mount of the calibration fiber of the top-calibration unit. To enable proper positioning of the calibration fiber in front of the pivot point of the arm. X-pos 2.75mm, Y-pos 3.15mm.

    Enlarged the 4 screwholes of the adapter plate holding the FIES fibers. This to allow the plate to be shifted to the left. No shift applied yet.

  • 25 Nov 2009: swapped to bundle C .

  • 26 Nov 2009: moved the adapter plate 2mm to the left to allow F1 to be in the FOV of Stancam direct imaging.

    Aligned fiberhead with telescope using a laser mounted in the adapter V-block. Tuned the rotational alignment slot, and the two rear support ball-screws of the head.

    With bundle C repositioned the XY slide to get maximum counts: X-pos 2.8mm, Y-pos 2.8mm. Calibrated the arm positions for fib8server.

  • 17 Feb 2010: lightning; power cut. Verushka down.

  • 19 Feb 2010: broken power supply of verushka was replaced

  • 22 Feb 2010: FIES on pump; connected pressure sensor

  • 23 Feb 2010: FIES off pump; disconnected pressure sensor

  • 5 Mar 2010: CCD reference temperature set form -120 to -110, because of cryotiger problems

  • 8 Mar 2010: shielded CCD-controler display as it caused lightleak in exposures with open shutter (fault 1312).

  • 3 Apr 2010: CCD reference temperature set back to -120 Centigrade

  • 12 May 2010: installed ADC in telescope adapter

  • 19 May 2010: took out the FIES pickoff mirror from te adapter, to measure reflectivity, and to use as a model for glueing a mounting T-piece to a new mirror. Old mirror put back, and aligned with laser mounted in fiber-head V-block.

    On-sky tests showed that subsequently the StanCam pickoff mirror was misaligned by about 1 arcminute.

  • 22-24 May 2010: powercycled cryotiger (22/4), powered off tiger and pumped cryo (24/5)

  • 26 May 2010: replaced the old FIES pickoff mirror with the new shiny mirror from Edmund, with enhanced aluminium coating. The new mirror has better reflectivity in BVR by a factor of 1.1

  • early June 2010: new small stones (gravilla) on roof of building, in order to better protect against excess solar heating

  • 11 June 2010: new CCD pressure sensor installed, electrically isolated from CCD controller

  • 14 June 2010: reglued STANCAM pickoff mirror in situ

  • 18 June 2010: external lights on CCD controller and pressure sensor blocked, as light leaks were reported

  • 22 June 2010: disconnected pressure sensor because it generates light, which goes through the tubes into the dewar

  • 23 June 2010: reglued STANCAM pickoff mirror on bench. Results look OK, rotator center quite centered

    Fixed ADC prism drive

  • 24 June 2010: installed light shielding in pressure-sensor tubing: no more stray light on CCD

  • 30 Aug 2010: changed top-unit ThAr lamp for a new one

  • 21 Nov 2010: cryo pumped just in case, as the pressure sensor reported weird readings

  • 23 Nov 2010: cryo warmed up by itself. Valve and sensor were dismounted and their connection parts were polished ; O-rings were cleaned and greassed; cryo pumped again

  • 25 Nov 2010: another spontaneous warmup. Replaced the FIES pressure sensor with another one taken from one of the pumps. Also the 'new' non-conductive mount of the pressure sensor was replaced with a metal one, as the 'new' mount was considered non-robust by Pedro; warmed-up and pumped

  • 17 Jan 2011: disconnected the cryo-pressure sensor, as the observers of 18/1 indicated that the extra noise may affect their data.

  • 21 Jan 2011: 20:00 cryotiger off 10 minutes and on again, connected pressure sensor. Ref. temp to -110

  • 22 Jan 2011: pressure sensor disconnected; FIES detector ref. temp. back to -120

  • 3 Feb 2011: dismounted the FIES dewar for maintenance. It was suspected that the vacuum seals were not made to precision, possibly leading to vacuum problems. Disassembled front flange, window and ring, sensor and valve mounts and right-angle pipes. Some scratches under Orings were polished, cleaned with acetone and baked in the oven for ~1 hour at ~80 deg.

    Solved the vacuum connections for the valve and sensor, now they cannot turn and the sensor is galvanically isolated, the sensor is back to original position but without the 10cm extension. We should try to keep the vacuum clamps as they are now because both have been adapted. The 2 right anlge bits got new O-rings in the dewar side.

  • 4 Feb 2011: Started to pump already in the sprctrograph room with evrything in position, connected and closed, at ~13 hours. The ref temp was set to 65 degrees to bake the baby.

  • 5 Feb 2011: at ~9:00 started cryotiger reset ccd ref temp, cover the sensor led and cover temp controller display.
    Pump off and out of the spectrograph room at ~15:15 hours.

  • 6 Feb 2011: Switched off cryotiger for few minutes to see if any reaction after powering on again, at ~9:05.

  • 8-11 Feb 2011: Changed broken focus-drive init switch, and encoder electronics. Refocussed spectrograph, with uninitialised but working encoder.

  • 3 March 2011: FIES cryo lost vacuum from ~1.2E-5 to ~5E-3 resulting in the increase in dewar and detector temperatures.
    Stopped the cryotiger, pumped, and filled the Cryotiger hoses from 215 to 247 psi.

  • 11 March 2011: lightnincg strikes.
    In the Fies building two RCCBs had tripped resulting in the CCD dewar warming up. The dewar was put on the pump and the next morning the CryoTiger was started and the pump removed. The FIES CCD and shutter seem to work. Bottom and Top calib units verified to work.

  • 8 April 2011: changed the top-calib ThAr lamp

  • 21 April 2011: low top-calib halogen light levels. GC checked lamp: seems OK

  • 12 May 2011: because of low Top-Calib halogen light levels checked the alignment of the calibration fiber where attached to the arm. After securing the rotational alignment (telephone number on blue-plastic part of plug should be facing DOWN) and fixing the rotational freedom of the calib fiber with fresh tape, both lamps give acceptable count rates. Both give expotimes less than 3 seconds for all 4 fibers, when using easythar/halo.

  • 4 June 2011: installed the new roof shade

  • 15 June 2011: replaced the stuck FIES main shutter by spare. Installed temporary cooling fan in the front room, as the current one was broken.

  • 27 July 2011: installed the protection for the exposure meter

  • 6 Sept 2011: pumped cryo; a long warm pump

  • 12 Sept 2011: pumped cryo; a short cold pump with cryotiger switched off

  • 4 Oct 2011: pumped cryo; a short cold pump with cryotiger switched off

  • 5 Oct 2011: pumped cryo; a long warm pump

  • 16 Oct 2011: CP: changed the vacuum valve and sensor. The sensor is the same but cleaned inside and more or less calibrated with the pump one, connected one in front of the other with a T bit. In the afternoon, the pump was started and after 10 minutes both sensor readings were close to 10-5.

  • 17 Oct 2011: CP: this morning by 10 hrs, the cooler was started. 3 hrs later, and since it was looking bad from the temperature point of view, worse than -160, it was applied Cox experience in stancam, power off 1/2 hour the cooler and back on. It got a bit better but still bad, so, 2nd try again by 15:30.

    During this 1/2 hour, I opened the cooler to make a drawing of the gas circuit and trying to understand the thing and where it can give problems in the way that after few days it can jump by itself to the proper temperature. The drawing was made and we can say that this is the same as a standard friger, or even like the turpol cooler. The unit was cleaned, especially the radiator. The only 2 possible places where we could get problems are a capilar tube, by getting a blockment, or a by-pass valve, which could get stuck. The rest is absolutely passive, tubes and a compressor. I was expecting the unit to fail again after powering on to give some hits in the named parts, but it did not fail. The unit is open now on the floor, I did not want to change any conditions since it shows correct values. Also because I want to try something if it fails again.

    Now we do not know what is really the thing, we know that the unit is working close to a limit which we do not understand. What was the solution? Twice powering off or the cleaning?

  • 31 Oct 2011, CP: I´ve installed a new O-ring for the front window, 60X3 after washing it with alcohol and add some vacuum compound. Also, I´ve changed the sensor heater one too. Now the cryostate is back in place, cooling hoses connected, spectrograph white insulation mounted back and the vacuum pump doing the work. After 1/2 hour, more or less, the pump sensor shows 6.5e-5 and cryostate sensor, read from the controller, 1.2e-4.

    The following morning, pump pressure 4.8e-6, controller pressure 2.9e-5.

    Just for your info, before installing the front window, I´ve polished the main front plate and the small holding plate in the O-ring sits using the small lathe in workshop, now they look clean and shinning. Also, by hand, the small block opposite to the cool head. All O-rings were washed, and greased with the vacuum compound.

  • 7 Nov 2011, CP: Adjusted CCD-cable positioning as the read-out noise was bad.

  • 8 Nov 2011, CP: closed the cryotiger box that was opened 17 Oct.

  • 19 Nov 2011, GC : The pump was put on, ran for over a hour before opening the valve. The CryoTiger was switched off for approximately 1/2 hour. The temperature rose quite fast, quicker than on previous occasions and some lose of vacuum occurred. When the CryoTiger was switched on the temperature initially fell reasonable fast but stopped at -185C and started to rise, the pressure also followed. The pump was left on, when the vacuum reached E-5mbars the valve was closed. The expected increase in pressure occurred but the temperature didn't follow. After waiting over half hour to see what would happen the instrument was closed so if the weather cleared it could be available for the end of the night when it would have been required.

  • 21 Nov 2011: switched off cryotiger, put on pump

  • 24 Nov 2011, GC : the Moxa box (Nport 5230) that communicates with all the temperature sensors (except the two web sensors) was replaced.

  • 24 Nov 2011, CP : replaced cryotiger compressor by spare; hoses and coolhead are still the same. Stop pumping at 14:45.

    The front-room temperature lowered down 2 degrees.

  • 29 Jun, 2012: PB exchanged top-calib ThAr lamp

  • 2-3 July, 2012, CP: cryo cycle: warmed-up, pumped, and changed faulty pressure sensor

  • 18-19 July, 2012: GC fixed a loose wire in the top-calib electronics box

  • 17-26 July, 2012: CP+TA tried to tackle the increasing room temperatures. Tried opening the doors to the front room, with and without ventilator positioned on the top of the white box. RC installed a second fan inside FIES room 20/7. GC switched off the small fan 26/7.

    Setting the cooling in the front room lower and opening the doors to the FIES room did not have any effect, but installing a small ventilator on top of the white box and aimed at the door did cause the FIES room temperature to drop. The temperature measurements in the FIES room (Aarhus 1, 2 and 3) stabilised at values slightly higher than under normal circumstances. When the outside temperature started to decrease the ventilator was switched off and the door between the FIES and front room closed.

  • 5 Aug 2012, Another warm period: switched on the ventilator at low speed, with only one door to the front room open.

    Found that the fan was installed with an extension cord with a bright red switch light. I removed the extension cord from the room.

    Re-taped green LED of CCD pressure sensor, as stray light was recorded on the chip in faint-star exposures.

  • 6 Aug 2012, GB Switched off the fan and closed all doors.

  • 13 Aug 2012, GC switched on fan and opened 1 door. Later switched off and closed by TP.

  • 14 Aug 2012, TP switched on fan and opened 1 door.

  • 26 Aug 2012, switched off fan and closed door.

  • 3 Sept 2012, JC changed the network switch in the FIES building to a more modern one. In the past it has been suspected of having ports that are bad.

  • 3-4 Sept 2012, GC worked on faulty temperature sensors, and replaced corresponding MOXA unit. One sensor control unit out of order.

  • 18 Sept 2012, JC changed the adapter of the fiber-to-wire network connection of the FIES building.

  • 19 Dec 2012. Removed tiny piece of obstructing glass from calibration ARM pivot.

  • 26 Dec 2012. GC installed new fire extinguishers in the FIES building

  • 3 Jan 2013. GC installed new top-calib ThAr lamp.

  • 12 Jan 2013. CP installed new bottom-calib ThAr lamp.

  • 26 April 2013. JJ, TK had problem with power socket in front room. Entered main room for access to main fuses.

  • 29 April 2013. GC fixed power socket, and switched of the heater in the front room.

  • 15 Aug 2013. Hot-period: TK opened the door to the main room and installed fan pointing at the front room. Even though the temperatures seemed cooler in the front room, the main-room temp went up considerably after this opening of the door.

  • 17 Aug 2013. Hot-period: PS aired the building in the morning, left the door to the main room open.

  • 18 Aug 2013. Hot-period: GC closed the door to the main room. Air temperature in the main room is back to normal, but the FIES temps are still high.

  • 21 Aug 2013. CP aired the building for a short while, and switched off the fan that was still on top of FIES blowing at a closed door to the front room. The past 3 days the fan caused a different (probably more homogeneous) vertical temperature profile in the FIES room, which resulted in relatively high temperatures under the table and inside FIES (0.5 degree warm). The ceiling temperature was normal. After switching off the fan the FIES temperatures are slowly decreasing toward their normal values.

  • 19 Sep 2013. CP moved the cryotiger from under the verushka computer table to the opposite corner, close to the entrance of the building.

    We note that the dewar temperature gradually but only slightly warmed up starting the day of the move. This triggered a temperature alarm 14 Oct, and we decided to pump (see next items).

  • 7 Oct 2013. CP and a construction guy made a hole between the front room and the FIES room to allow vacuum piping to be installed, for the future pumping of the cryostat with the pump located in the front room. This to avoid temperature fluctuations while pumping, as the pump heats up the FIES room.

  • 14 Oct 2013. As the dewar pressure more than doubled the value after the previous pumping, 14 months before, we decided to pump.

  • 15 Oct 2013. GC: After sitching back on the cryotiger, it suddenly failed to cool after 10 minutes. The fans on the compressor were working OK. The flow was reversed, to clean a possible blockage, which then caused the system to resume cooling. As one of the tubes has a broken auto-seal, some gas escaped. After the hoses were reconnected in the normal direction of flow, the system was topped up with coolant.

  • 14 Nov 2013. General power cut with generator failure. GC writes: "FIES CCD has warmed up and Ilya doesn't necessarily need it tonight. I propose letting it finish warming up, I put the pump on tonight and it will be ready by tomorrow night."

  • 15 Nov 2013. FIES back in operation after warming up and pumping.

  • 17 Dec 2013. Changed bottom ThAr lamp.

  • 6 Feb 2014. GC installed some terminators in the top-calib electronics box, and installed a separate power supply for the MOXA boxes.

  • 12 May 2014: the dewar vacuum was steadily rising, and a pumping cycle was planned, also because the cryoTiger needs to be refurbished to allow the filter to be refreshed. GC+PB swapped the cryoTiger compressor for a brand new one. They also replaced a damaged supply hose.

    It took some time and effort to get the new cryoTiger going, some flushing of the coolhead was done. Additionally, the vaccuum ducting of the pump was failing. The cryoTiger was finally going on 16/5.

  • Apparently, the above introduced condensation on the CCD. Similar to what was seen in the flatfields obtained after the Oct 2013 pumping, a much stronger pattern in the extracted flats was evident. The pattern does not change when looking at flats from day to day.

    A new warm pumping cycle was executed 26-28 May, after which the pattern was gone.

  • 5 June 2014. CP installed a new air cooling system (a heat pump) in the FIES front room that is decoupled from the general telescope cooling system.

  • 20 Nov 2014. Lightning strike. Some systems affected (?): notably a MOXA box connectivity problem (temp sensors) got repaired by the lightning.

  • 9 Feb 2015. New Photron lamp installed in the top calibration unit. This is the first Photron lamp we used. Some Hereus and 2 Photron lamps remain for spares. It is likely that the replaced Hereus lamp is not broken and still can be used.

  • 16 Feb 2015. GC fixed the non-reporting temperature sensors in the FIES room: the problem seems to have been a dodgy connection inside the FIES instrument room; for a change it was not the MOXA-box problem.

  • 10 Apr 2015. The temperature sensors reported to the WWW for less than 2 weeks following the latest repair.
    GC : as a temporary working solution I've bypassed a pair of the temperature reading modules in the FIES building so you now get sensor reading from s531, s532 and s533 (the so-called Aarhus sensors).

  • 18 May 2015. GC changed the temp-sensor wiring in the FIES building such that only s531, s532 and s533 are functioning, rather than none.

  • 20 May 2015. We decided to pump as the pressure was gradually getting higher. No panic, just a precaution for the Summer when staff will be on holiday.

    GC: as usual thing never go as planned! Switching on the CryoTiger the temperature started to drop but after a short while started to rise again, bummer!!! It was noted that the working gas pressure was virtually zero (classic case of blocked cold head) so decided to switch the unit off for ~20 mins and restart it. Again a small drop in temperature then increasing, so didn't help.
    Since I have no intentions of playing with the gas hoses and the inflammable gas on my own I waited for Pere to arrive, then swapped the return and supply hoses at the Compressor end.
    When the CryoTiger was started it could be seen that the working pressure went to ~25 psi, which is reasonable and things started to cool.
    Hoses swapped back after ~ 1 hour (13:00h). This time the working pressure showed around 30 psi (which is normal) and cooling.
    Note the static gas pressure is ~255 psi, which is very good.

    JHT: i checked the RON, focus, and for light-leaks: all good!

    JHT: but unfortunately the fixed-pattern has re-appeared in the flatfields.

  • 22 June 2015: as of June 2, the top ThAr expo times have increased by a factor 2.5-3 for fibers F3 and F4. I re-aligned the calibration fiber position using the XY slide. The ThAr expotimes, for a PHOTRON lamp, are now 3 or 4 sec for all 4 fibers. I had to compromise the Halogen throughput to achieve this: between 2 and 28 sec for the 4 fibers.

  • 24 June 2015: after updating the fib8server software ARM positions, all easythar/halo expotimes are quite optimal again.

  • 17 Sep 2015, three jumps of 2 degrees occurred in the front-room from Aug-Sept. The August jumps possibly related to the big power outage.

    GC updated/resetted the settings of the heat-pump: "I've changed the FIES front room set-point to 19C and replaced the batteries in the controller (though probably not required). Also the controller was set to cooling only and maximum fan speed, I've set it to automatic such that it will maintain the temperature (cooling or heating when necessary) and put the fan speed to automatic."

  • 22 Sep 2015: Pepe Frio's team replaced the external coolant valve control unit. The heat-pump/inverter set-point has been set back to 21C. Seems to be working again.

  • 10 Dec 2015: For the latest ~10 days the dewar temp is rising much faster than normal (0.4 degrees in 10 days), while the CCD vacuum rises at its normal slow rate. It was suspected that the cryotiger lost some of its performance. The working pressure of the CT gas was 21 PSI, the static pressure around 215 psi.

    After trying the effect, found non beneficial, of power cycling the CT, it was decided to top up the CT with gas from a bottle with only 18 psi on its gauge. After filling, the bottle still got 10 psi to go. The CT went up to 255 psi.

    This was done while pumping: it took a long time until the pump pressure was below the dewar pressure in order for the main valve to be opened. The CCD pressure before pumping was 22 pressure units, and after it was 31 pressure units (0.0000031 on the TMS).

  • 14,15 Dec 2015: CP: In this order: pumped with cooling off from day before yesterday afternoon until yesterday morning; started the cooling with reversed flow for 1 hour; stopped cooler, put the hoses back to normal, and switched-on the cooler again; stopped pump and removed it from the spectrograph.

  • March/April 2016: new CCD. Initially the spare cryotiger, with a new filter installed, gave severe cooling problems. An attempt was made to use the old-CCD's properly working cryotiger. The first attempt only shortly on 1/4. As a consequence of the second attempt the old CCD warmed up and had to be repumped. It was decided to keep the CCD at +40 degrees centigrades while warm, and this resulted in the disappearance of the long-standing condensation pattern seen in flat fields. This warm period lasted from 7-11 April, while at +40 from 9-11 April.

  • 6 June 2016: JT+GC+PB: swapped the broken bifurcated calibration fiber (top calib unit) for the spare. The fiber-coupler thread of the arm-pivot is worn. The custom-made piece may need to be remanufactured.

    The XY-slide positions and software arm positions were optimised, the first rather hastily. Exposure times are about a factor 2 longer, except for F3,F4 ThAr.

  • 10 June 2016: the 15V power supply of the StanCam controller blew up causing loss of StanCam. See fault DB 2507, 2528

  • 25 June 2016: In the afternoon readjusted the FIES calibration fiber alignment. With new fiber connector piece done by Pedro.

  • 21 July 2016: Hot room. Start of warm summer weather. On 1/8 the fan was switched on with the intermediate door open.

  • 28 Aug 2016: leaks in the cryotiger hoses of StanCam

  • 22 -30 Sep 2016: installation of new FIES CCD15 with new controller. We had to make the mounting flange of the new CCD fit, by mechanically taking 2mm off the diameter.

    New ducting for the pumping from the front room was finalised. New ducting for the exhaust of the controller was installed. All this required new holes to be cut in the white box, while old holes were closed.

    The new ccd3-fies computer was installed in the front room, and a light-weight monitor was installed to replace the old CRT monitor for verushka.

    The cross-disperser was tweaked by 5 full turns (clockwise) of the adjustment knob, in order to shift the wavelengths sampled by the new 10% larger CCD towards the red. FIES was refocussed by only 20-25 FIES focus-drive units.

    New bright LEDS (CCD controller and pressure sensor) were taped off, but stray undispersed light hits the detector (at high X values for amp B) when the shutter is open. I placed the rubber mat over the shutter and fiber-slit parts, to get rid of most of the stray light.

    The old CCD13 was left on a small table in the FIES room, while connected to get the telemetry values on the WWW. With its CryoTiger switched on.

  • 16 Nov 2016: JC installed a switch on the CCD controller that can be operated over ethernet.

  • 9 Dec 2016: unintentional warm up of CCD15: pumped cold

  • 17 Jan 2017: unintentional warm up of CCD15: pumped warm with CCD-heater set to +30

  • 7 Mar 2017: GC et al. installed: a cleaned pressure sensor on the CCD cryo as the old could no longer measure below 0.004mBar. The old got cleaned and is now the spare. The cryo was pumped warm without the CCD heater switched on.

    Furthermore, the CryoTiger gas-pressure sensor (WWW reporting) was mounted, and a new extra filter at the cryostat end. The WWW reporting has yet to be activated.

    The front room was tidied upto make room for the new big Lakeshore temperature loop controller. As part of this the CCD15 data fiber was rerouted towards the ccd3-fies computer over the cable ducts.

  • 8 Mar 2017: the ccd3-fies computer needed rebooting, as it crashed during readouts. It was found that the ccd3 PC interface board and the its PCI slot in the PC both do not function anylonger. A spare board was installed in an adjacent PCI slot.

  • 29 May, 5 and 9 June 2017: CP and PB prepared for installation of active cooling of the front room, by installing holes in the doors between the two rooms for inlet and a fan, and ducting. The cooled front-room air will be used to refresh the main room when it is too hot.

  • 14 June 2017, Tapio and i taped and closed the wide slots between the white box and the bottom of the optical table. And we disabled the fan in the white bix that sucks 'cool' room air into the spectrograph. This was done in order to avoid dust coming into the spectrograph when the new room-to-room cooling fan will be active. This effectively raised the temperature inside the white and black box by 0.5 degrees.

  • 19 June 2017, with PB and CP removed spare bundle B from the telescope. The glass slit was damaged and the bundle cannot be reused without repair.

  • 30 June - 8 July 2017. Installation of the new fiber bundle D. With Julian Stuermer and Andreas Seifahrt.

    Unfortunately, during the installation process, the med-res fiber F3 of bundle C ceased to function properly.

    The new bundle was aligned with the telescope, firstly using a laser mounted in the tip-tilt stage of the fiberhead, and secondly by observing the FRD in the spectrograph room while observing alpha Boo. The FRD was measured using a photographers camera, setup to monitor the size of the projected beam, while adjusting the tip-tilt stage.

    The high-res fiber F4 of the new bundle was aligned with the grating, which gave good throughput results also for the low-res fiber F1. The fiber-exit lens needs to be redesigned to get the beams of the med-res fibers properly aligned with the grating. Currently they give very poor throughput.

    The new telescope coupling came with a new calibration arm and a new fiber mask mechanism, both with encoders causing stray light in the telescope adapter in the i and z bands. This is problematic for StanCam in direct imaging mode. Unfortunately the encoders cannot be switched off in software, in a way allowing reliable operation of the mechanisms after switching them on again.

    The new bundle was left installed for scientific use.

  • 10 August 2017, with Pedro changed the mounting of the top ThAr to get more light into the new ThAr calibration fiber. We made 2 new sleeves for the ThAr lamp (called "fatning" in the drawings). One with inner diameter of 38.0 mm and one with inner diameter 37.15 mm. The second allowed for good mechanical stability of the lamp. Pedro made a new clamping screw in the hull housing the lamp, such that the lamp can be clamped, and no tape is necessary for that purpose anylonger.
    See the original drawings.

    The new sleeve allows the lamp to be moved into the housing, and we found that highest counts are achieved when moving the lamp approximately 28mm towards the lens and fiber, with respect to its original position. Finally the sleeve was adjusted with an extra spacer ring such that highest counts are achieved when pushing the lamp fully into the housing. I estimate that we get on the order of 50% more light into the fiber.

    When changing the lamp, care must be taken to first unscrew the new clamping screw. I should have taken a picture of it ....

  • 26 October 2017, the ADC and the FIES pickoff mirror were both cleaned (AD)

  • 3 Jan 2018, after many TMS warning emails, the FIES cryotiger was power cycled 3/1/2018 by Sune, which cleared a 'blockage' and fixed all readings to normality.

  • 5 Jan 2018, GC installed a new ThAr lamp in the top unit. The old one seems not broken and is likely still useful.

  • 22 Feb 2018, CP+PB flushed the coldhead of the CCD cryostat, and installed a new cooling-gas-flow filter to replace the home-grown filter near the CCD. The cryostat was pumped with CCD ref temp set to +30.

  • 28 Feb 2018, GC: another warm pumping cycle induced by the severe lightning strike of 25 Feb.

  • 29 Feb -- 6 Mar 2018, Various fixes related to the lightning strike; both front-room cooling systems were fixed; replaced the blown Moxa box ( in the FIES front room that monitors the Aarhus sensors and CryoTiger gas pressure gauge. FIES and StanCam cooling-gas pressure gauges both fail after the lightning.

  • 21 March 2018, the broken power supply of verushka was replaced. It broke due to another power outage.

  • 17 April 2018, tested a new calibration fiber for throughput; not any better.

  • 19 April 2018, GC fixed computer communication issue in front room; probably pending since lightning strike of 25/2/2018. The problem caused non of the calibration-unit mechanisms and focus drive to work.

  • 30 April 2018, swapped the new bundle D spectrograph-coupling for a new version that is to be used together with the original long fiber-exit lens. This to try to make the med-res fibers usable again.

    FITSheaders refer to bundle D2 as of May 15.

  • 10 May 2018, refocussed the new bundle D2 spectrograph-coupling. There was severe astigmatism for the fiber-exit-train position used since 30/4/2018. The fiber-exit-train was moved 8mm towards the spectrograph, in order to get good focus in both directions. This is compensated by a negative change of the focus drive by about 20-25 units (very imprecise).

  • 5 July 2018, opened the spectrograph and grating enclosure to take more pictures and measure some items, for Niels Christian Jessen, in relation to a possible grating-tank upgrade.

Back to top Last modified: 01-Aug-2018